OpenStudy (anonymous):

What are all the real zeros of y=(x-12)^3 - 10?

OpenStudy (anonymous):

\[(x-12)^3-10=0\]add 10 to both sides \[(x-12)^3=10\]take the cubed root \[x-12=\sqrt[3]{10}\] add 12 to both sides \[x=12+\sqrt[3]{10}\]